### v1

# BlockLLM: Multi-tenant Finer-grained Serving for Large Language Models

Abstract
The growing demand for Large Language Models (LLMs) across diverse applications has prompted a paradigm shift in the design of deep learning serving systems. Deploying LLMs, especially in multi-tenant environments, presents considerable challenges due to their high computational and memory demands. We present BlockLLM, a serving system that exploits the potential of sharing components among fine-tuned LLM models to offer an efficient and flexible solution for LLM workloads. BlockLLM partitions the models into finer-grained blocks to enable the reuse of model components and independent provisioning to improve the computation efficiency. BlockLLM consists of an offline block zoo, for storing the blocks, and an online system to serve the requests through chains of blocks. It offers multi-fold flexibility: (1) Adaptive assembly of block chains on-the-fly is achieved with the help of equivalence evaluation among blocks in the zoo. (2) We enable per-block batch size and configure best-effort KV cache coordination at individual block level. (3) We adopt speculative execution and locality-aware block placement to mitigate the communication costs from dynamic block resource allocation. Our evaluation demonstrates that BlockLLM reduces memory and storage footprints and improves computation efficiency, outperforming existing serving approach in 95%ile latency and GPU utilization by 33.5% and 20.1%, respectively.
1 Introduction
The rise of Large Language Models (LLMs) marks a transformative milestone in deep learning. Their unprecedented abilities in natural language processing [28], from language translation [6] to question-answering [1] and complex reasoning [20, 40], are reshaping technological frontiers. Yet, the strength of LLMs—rooted in their vast parameter spaces—comes at the price of significant computation costs [5]. The deployment of LLMs is an arduous endeavor, often necessitating the use of hundreds or even thousands of extremely expensive computation devices like GPUs, thus imposing great challenges on unleashing LLM’s full potential for more people [36, 3].
With the growing reliance on LLMs, fine-tuning has emerged as a critical technique to efficiently adapt foundation models to handle various tasks across different domains [16, 9]. Fine-tuning updates existing parameters or introduces additional parameters using domain-specific data. As LLMs become ubiquitous, serving these fine-tuned models from a multi-tenant cluster, whether in a private cloud for first-party services or a public cloud for third-party users, is starting to emerge as an important consideration.
Recent serving systems and approaches focus on model-internal optimization, such as the acceleration of matrix computations [8, 2] and minimizing device memory fragmentation [18] to enable parallel decoding to circumvent the limitations imposed by auto-regressive models that process a single token at a time [11, 26]. While these advancements significantly accelerate the inference process for individual LLMs, the challenge remains to serve multiple tenants’ fine-tuned models in a shared cluster. Each LLM requires substantial resources, and the hardware cost makes it prohibitive to liberally expand the cluster to add dedicated resources for each tenant. Thus, the question is: how can LLM serving meet the latency performance goals of tenants’ fine-tuned models while also ensuring high throughput and optimal utilization cluster-wide?
We introduce BlockLLM, an efficient and flexible serving system tailored for LLM workloads in a multi-tenant cluster. BlockLLM enables LLM model component reuse by breaking down an LLM into smaller and shareable components which we call “blocks”. The central observation forming the basis for BlockLLM is that LLM fine-tuning offers the opportunity for sharing. Parameter-efficient fine-tuned models have an inherent modular architecture where parameters from specific layers are added or altered. Our observations with full-parameter fine-tuned models indicate that certain model components produce outputs with high similarity, suggesting their potential for sharing as well. Block sharing reduces both memory and storage demands. Furthermore, by reallocating memory - previously consumed by redundant parameters - toward supporting larger data batch size, BlockLLM achieves higher computation efficiency and overall throughput.
Figure 1: System architecture. $b_{x}$ are blocks. ‘R’ are requests. $SB_{3}$ is the surrogates of Block $b_{3}$.
BlockLLM consists of an offline “block zoo”—a repository of LLM blocks, and an online serving system capable of dynamically scheduling blocks to serve various applications (Figure 1). BlockLLM’s practice brings three opportunities. (1) BlockLLM can support adaptive serving, which enables on-the-fly assembly of blocks for each request, moving beyond the traditional static chain of blocks. (2) Each block can be individually configured, allowing for customized batch sizes and autonomous handling of queued requests. (3) BlockLLM enables dynamic resource allocation, with each block capable of independently scaling and being placed without the constraints of model-level boundaries.
Several challenges arise in realizing BlockLLM with full flexibility. First, the feasibility of adaptive serving hinges on an in-depth understanding of the equivalence among blocks. This necessitates a mechanism to establish inter-block relationship within the block zoo. Moreover, LLMs store intermediate results in KV cache [31] to avoid recalculation due to their auto-regressive nature. The independent configuration of each block leads to overlapping lifecycles of different requests, demanding stateful coordination to maintain the coherence of the KV cache across concurrent requests. Lastly, resource allocation without model-level boundaries brings extra inter-block communication, thereby imposing additional latency costs that are not easily mitigated. In BlockLLM, we propose integrated solutions for these challenges.
First, to ascertain the equivalence of blocks within BlockLLM, we consider two methods subject to block architecture. For blocks of the same architecture that naturally support seamless request routing, cosine similarity metrics are utilized to measure parametric differences. For blocks of disparate embedding sizes, the challenge is more intricate. We compare the output distribution of vocabulary probabilities,
which serves as an indicator for functional equivalence between transformer layers. Inspired by the techniques of merging computer vision models [30], we design a generalizable stitching block as an intermediary to route requests among these blocks so that the chain of blocks to serve each request can be adaptively adjusted online.
Second, to manage the KV cache generated by different requests, BlockLLM implements a principled approach to dispatch requests. Per-block configuration introduces complexity to the handling of the KV cache, particularly as requests might be allocated to devices that do not currently hold the corresponding cache, necessitating either cache communication or recalculation. We analyze the trade-off between the recalculation and new IO costs and use it to compute different priorities for different candidate block instances, including the blocks specified in the chain and its equivalent blocks. We adopt a best-effort KV coordination policy where devices possessing the KV cache are prioritized when dispatching the request.
Third, to address the inherent latency costs posed by inter-block communication, BlockLLM integrates speculative execution for identified bottleneck blocks, thereby processing potential request pathways ahead-of-time. We use surrogate models to predict the outputs of blocks, facilitating computation ahead of actual output. In parallel, we implement a locality-aware strategy for block placement, putting interdependent blocks nearby, preferably within the same server. This approach reduces the frequency of costly inter-server communications and leverages the high-speed intra-server connections to expedite data transfer and reduce inference latency.
We summarize our contributions as follows:
  • We leverage the LLM fine-tuning characteristic and demonstrate the effectiveness of finer-grained LLM serving in terms of reducing memory and storage requirements and optimizing computation resource utilization.
  • We construct a block zoo that stores LLMs in finer granularity to enable block reuse and establishes the relationship among these blocks to facilitate adaptive online serving.
  • We build an online serving system to improve cluster throughput. It strategically coordinates the KV cache by best-effort request dispatching and minimizes communication overhead with bottleneck-pinpointing speculative execution and locality-aware placement.
  • We implement and evaluate BlockLLM, a multi-tenant serving system. We evaluate BlockLLM in a 12-A100 cluster. BlockLLM achieves comparable median latency with traditional per-model provisioning and reduces the 95%ile latency by 33.5%. The overall average throughput is elevated by 1.71x. GPU SM efficiency is improved by 20.1%.
2 Background and Motivation
2.1 Background
LLM serving. LLM-based applications have grown to be one of the major workloads in GPU clusters. Given their colossal scale, the burden of serving these models has emerged as a critical issue. In the offline stage, large storage is required to store the model checkpoints and engines. In the online stage, numerous powerful computation units (e.g., GPU devices) with large memory spaces are necessary to load and execute the models [35, 31]. LLMs’ unique auto-regressive nature necessitates caching KV matrices of tokens on devices to avoid redundant recalculations [32]. Every request has its own KV cache, which progressively grows in length as the auto-regressive process continues. Moreover, when models are inevitably deployed in a distributed manner, a high-capacity network is indispensable to transfer the intermediate activations. Such a burden is particularly pronounced in multi-tenant clusters, which cater to diverse LLM applications, each with its dedicated models and performance objectives.
Foundation models and fine-tuning. LLM training follows transfer learning as a guiding principle. A foundation model is initially pre-trained on a broad spectrum of general data, which is then customized for specific applications through fine-tuning. This process involves introducing additional parameters or modifying existing ones within the model using task-specific datasets. Fine-tuning entails updating either the entire model, known as full-parameter fine-tuning (FF) [4], or altering a small subset of additional or modified parameters, referred to as parameter-efficient fine-tuning (PE) [25, 12]. PE examples include LoRA [17], BitFit [43], Adapter [13], and Prefix-Tuning [21].
Problem statement. The wide adoption of LLM necessitates an efficient multi-tenant serving system tailored for such workloads. Beyond fulfilling latency requirements, our goal is to elevate throughput and resource utilization in a cluster where diverse LLM applications are present.
2.2 Finer-grained Serving
Key idea. To enhance the efficiency of LLM serving systems, we advocate a finer-grained approach. Rather than treating each model as an indivisible unit, we propose BlockLLM, which partitions LLMs into smaller components provisioned independently. These components are envisioned as “blocks”. We could maintain only one copy of a block even if it is used by multiple models. The immediate benefit is the decreased demand for memory and storage.
Figure 2: Example of blocks for two models fine-tuned from the same foundation. We show one Transformer layer for simplicity.
PEFT models are particularly well-suited to this approach due to their inherent modular architecture. For example, as illustrated in Figure 2, two models could be broken down into five blocks, with blue blocks being common components shared between them. Moreover, our analysis of FPFT models indicates that blocks with varied parameters often produce highly similar outputs. Figure 3(a) presents the parameter cosine similarity of each Transformer layer between LLaMA and Vicuna/Alpaca (FF from LLaMA). The average similarity is 0.9927. This observation underpins the feasibility of reusing blocks across different FPFT models, even when parameter values differ. Enabling the reuse of model components would improve resource efficiency by letting device memory hold more critical data.
(a) LLaMA 7B and Vicuna 7B.
(b) LLaMA 7B and Alpaca 7B.
Figure 3: Cosine similarity of parameters in each Transformer layer of two models.
Opportunities. BlockLLM consists of an offline block zoo (4) and an online serving system (5). We enable block reuse in block zoo and establish connections among blocks. In the serving phase, BlockLLM deploys the blocks onto the cluster and assembles a chain of blocks to serve each request. Each block instance is provisioned independently. Block-granularity provision brings three significant opportunities for enhancing serving flexibility.
O1: Adaptive serving. Deployment and serving can be highly adaptive. Blocks allow for dynamic assembly into a variety of model configurations. Serving a request is not constrained to a predetermined chain of blocks. Instead, it can be dynamically adjusted in real-time, based on the deployment state, to improve efficiency. The key to realizing this fully adaptive serving hinges on accurately determining whether the blocks are functionally equivalent (4, 5.3).
O2: Per-block configuration. Each block can have its unique configuration. First, the batch size of each block could be individually configured. Intuitively, a block shared by multiple applications should have a larger batch size to efficiently handle the collective load. Second, each block maintains its own request queue and manages the request state independently, thus autonomously deciding how to process the requests. A critical consideration in this configuration is how auto-regressive models can be integrated into this paradigm, where the KV cache should be managed in a cohesive approach (5.1).
O3: Dynamic resource allocation. BlockLLM enhances resource allocation with its ability to independently scale each block. Blocks that are frequently accessed or computationally intensive can dynamically be allocated with more resources. It also liberates block placement from the constraints associated with monolithic model architectures. However, it does introduce a challenge in the form of increased communication costs. Effectively mitigating these costs is crucial to achieving truly versatile resource allocation within BlockLLM (5.2, 5.3).
2.3 Comparison with Existing Work
Existing serving systems. Existing serving systems for LLM workload follow the practice of serving traditional DNNs. Each model operates within a dedicated engine and is provisioned as a monolithic unit. An alternative approach is multi-tasking. A typical example is S-LoRA [34], which exploits memory paging to pack multiple LoRAs within one engine. Another line of optimizations focuses on the model internals—like accelerating Attention layer computations (FlashAttention [8]), managing KV cache (PagedAttention [18] in vLLM), and efficiently scheduling varying sequence lengths of requests (Orca [42]). These efforts are complementary to BlockLLM. BlockLLM tackles the inefficiency of serving LLM workloads in a multi-tenant cluster by leveraging the characteristics of LLM fine-tuning. Block-granularity provisioning improves the serving efficiency from two perspectives.
Model PE % of Shared Params.
LLaMA 7B LoRA (Green) 99.94
Adapter (Purple) 92.62
Prefix (Orange) 99.88
GPT-NeoX-20B LoRA 99.52
Adapter 92.52
Prompt 99.98
Table 1: Percentage of shared parameters of different PEFT techniques.
Figure 4: Architectural changes of PE.
Figure 5: Redundancy and switching overhead in the existing serving system. x×$\times$×y indicates we have x foundation models and y models fine-tuned from each foundation. Therefore, xy models in total.
Reduced memory and storage. Models fine-tuned from the same foundation model inherit a significant number of shared parameters. We detail various fine-tuning techniques and the extent of parameter modification in Table 1 and illustrate the architectural changes in Figure 4 with shared parameters depicted as gray boxes. Existing serving systems ignore this aspect, leading unfortunately to the redundant storage of identical model pieces. In Figure 5, we present the percentage of the model parameters being redundant. With more diverse applications served in a cluster, the size of redundancy is more significant. When the cluster hosts 15 LLM-based applications that are adapted from three foundation models (3rd bar), the 92.1% redundancy takes up to $\sim$147GB. In terms of memory-related operations, this redundancy becomes particularly instrumental, where models are frequently swapped in and out in response to requests from different applications. A considerable portion of the switching overhead (85.4% with 20 applications) is wastefully spent on unnecessary replacement of the same parameters. With BlockLLM, the switching overhead is reduced to at most 12.2% with 9 applications and 16.5% with 20 applications.
No. of Apps Per-model Provision BlockLLM
Latency Throughput Utilization Latency Throughput Utilization
3 5.4 6.6 64.4% 4.6 8.9 76.5%
6 7.6 8.2 75.1% 5.2 12.3 79.4%
9 14.4 7.2 76.6% 9.9 14.5 85.3%
12 17.6 6.9 74.4% 12.4 15.6 89.6%
Table 2: Comparison of average latency (mins/request), throughput (tokens/second), and GPU utilization (SM efficiency) when the number of applications increases. GPU utilization is measured until the last request is completed.
Improved computation efficiency. Serving each model within a standalone engine presupposes a complete execution cycle. As such, per-model provisioning prevents the sharing of computations among models, even when they contain identical components. It is inefficient in a multi-tenant cluster where many different applications are served. Table 2 shows the average latency, throughput, and GPU utilization of a small cluster of four GPUs where 40 requests from three applications are served. When using per-model provisioning, the throughput and GPU utilization are low because the batch size is small, and the devices hosting less popular applications are severely under-utilized. Alternatively, with BlockLLM, we could achieve a 34.8% higher throughput. The gain primarily stems from three aspects: (1) the reduced overhead by loading smaller model pieces, (2) higher computation efficiency due to larger batch sizes for shared blocks (O2), (3) effective scaling of popular blocks onto the idle computation resources originally reserved by less popular applications (O3). We also measure the performance improvement when the number of applications the cluster hosts grows from 3 to 12 and the number of requests grows proportionally. The throughput improvement of 12 applications is 1.91x of the throughput improvement when hosting three applications.
3 BlockLLM Design
We provide an overview of BlockLLM and explain its three design challenges.
3.1 System Overview
Architecture. Figure 1 depicts the system architecture of BlockLLM. In the preparatory offline phase, BlockLLM hosts a repository—termed “block zoo”—that houses the LLMs in blocks. It is equipped with an automated mechanism that partitions the models into blocks and ascertains the equivalence among existing blocks. Additionally, the block zoo integrates a profiler to record the performance metrics and trade-offs pertinent to each block. In the real-time online phase, BlockLLM has a scheduler that orchestrates resource allocation and placement of blocks and processes the requests. It strategically schedules blocks onto devices, denoted as “block instance”. A BlockLLM agent resides on every device in the cluster. It maintains and monitors the block instances and the associated request queues. It is responsible for handling the requests, including managing the KV cache and transferring outputs to other block instances.
Workflow. During serving, the workflow is as follows. When a new request arrives, the scheduler assigns a chain of blocks based on the application it belongs to. The scheduler initially determines whether there is an available instance of the first block in the chain. If available, the request is forwarded to the target block instance. Otherwise, the scheduler either identifies a suitable alternative among the existing equivalent block instances or deploys a new block instance on an available device. Upon receipt of the request, the device’s agent directs it to the designated block instance for processing. Then, once the execution is completed, the agent is responsible for directing the request through the subsequent block in the chain. The process ends with the completion of the request signified by an EOS token, whereupon the final agent relays the output back to the scheduler, thus concluding the response cycle of the request.
3.2 Design Challenges
We explain three design challenges in BlockLLM, which are briefly mentioned in 2.2.
C1: Evaluate different-sized block equivalence. To achieve adaptive serving, establishing the equivalence between blocks is necessary so that chains of blocks for each request can be adjusted on the fly. Block-granularity model zoo offers improved manageability compared to full models due to their significantly smaller parameter size. Nonetheless, establishing such equivalence remains a complex issue. Despite the widespread adoption of the Transformer architecture in most existing LLMs, they often vary in size-related parameters like the embedding size. While metrics such as cosine similarity or L2 norm can be employed to assess the equivalence between blocks of identical architecture, evaluating blocks of differing architectures for equivalence is far from straightforward. Furthermore, even with an understanding of their similarities, it is not feasible to directly route a request to a block that differs in embedding size.
(a) Per-model provision.
(b) Block provision.
Figure 6: Illustration of coordinating KV cache when using block-granularity provisioning
C2: Coordinate stateful KV cache. In BlockLLM, sophisticated stateful serving is indispensable for auto-regressive LLMs. Existing serving systems operate by dedicating each model instance to process a single batch of requests until completion. This approach allows GPU devices to retain a singular, specific KV cache for the batch they are processing. However, BlockLLM introduces a more complex scenario where each block is configured independently. Each block instance may concurrently process multiple request batches over time, each generating its distinct KV cache. This interleaved processing approach, coupled with the auto-regressive nature of the tasks, implies that a request batch requiring subsequent iterations needs access to its original KV cache. However, there’s no guarantee that the same device that holds the KV cache will be available to process additional iterations of that batch. Figure 6 illustrates this challenge, showcasing how the blue request could encounter the problem of no available KV cache if dispatched to Block instance A2 under BlockLLM’s provision choice.
Figure 7: Components of inference latency of generating one token using LLaMA 7B in BlockLLM.
C3: Mitigate communication overhead. Versatile resource allocation of finer-grained blocks provides several benefits for online serving, as mentioned in 2.2. Yet BlockLLM incurs additional non-negligible communication overhead. Specifically, each time a request batch is directed to a different device, it necessitates intra-server or inter-server communication. This communication might act as a bottleneck for the batch, directly contributing to increased latency. The communication overhead also linearly grows with the number of block instances passed by a request batch during its lifecycle. Furthermore, the volume of data transferred escalates with the use of larger batch sizes per block. Figure 7 breaks down the latency components involved in generating one new token across various request batches. Without any optimization to the communication, for a request batch of size 32 and length 512, the actual computation constitutes 62.9% of the total latency, whereas the inter-server communication accounts for 36.4%.
4 Block Zoo
In this section, we mainly answer three questions:
  • How to evaluate equivalence among model components? (C1, 4.1)
  • How to partition LLMs into blocks? (4.2)
  • How to handle request routing between two equivalent blocks with different embedding sizes? (C1, 4.3)
4.1 Equivalence
PE models inherently possess equivalent parameters for reusing. We thus focus on verifying if equivalence exists among FF model components. Our goal is to find such equivalence so that some of the FF model components can be shared as well once they are partitioned into blocks. We consider two scenarios here.
Figure 8: Routing requests when equivalence is found.
Figure 9: Stitch blocks in different embedding sizes.
Models with an identical architecture. First, when two models share an identical architecture, we can establish the equivalence to optimize request routing. For instance, if tail layers (Layer $i$ to $L$) are determined to be equivalent, as depicted in Figure 9, we have the flexibility to route requests from head layers of Model B to tail layers of Model A, which becomes particularly useful when the tail layers of Model B is experiencing high load. This strategy is also beneficial in conserving resources and reducing loading overhead when no instance of the tail layer of Model B is actively deployed.
In this case, the degree of equivalence between model components can be systematically quantified. In BlockLLM, we employ the canonical metric, cosine similarity, to evaluate the parametric similarity of each Transformer layer. The similarity is defined as the weighted average of cosine similarities of all constituent parameters within each Transformer layer. That is, for layer $A_{i}$ and $B_{i}$ is:
$\vspace{-1mm}Eq(A_{i},B_{i})=\sum_{p=1}^{n}(s(A_{i}^{p})\times cos(A_{i}^{p},B_{i}^{p}))/\sum_{p=1}^{n}s(A_{i}^{p}),\vspace{-1mm}$
where $n$ is the number of parameters in $A_{i}$, $s(A_{i}^{p})$ denotes the number of values in the $p$-th parameter in $A_{i}$, and $cos(A_{i}^{p},B_{i}^{p})$ is the cosine similarity between the $p$-th parameters of $A_{i}$ and $B_{i}$. A higher value indicates a high similarity or equivalence between two Transformer layers.
Models with different embedding sizes. Checking equivalence among model components with different embedding sizes offers more flexibility in the sense that requests can potentially be routed to a smaller model component. It helps in balancing load and reducing latency, as processing through a smaller yet equivalent model component is faster.
Since size differences are common for models adapted from different foundation architectures, we focus on foundation model parameters. For example, the embedding sizes of 7B and 13B LLaMA are 4096 and 5120, respectively. Parametric cosine similarity is no longer applicable here anymore. Resorting to methods that enforce uniformity of dimensions, such as dimensional reduction, is impractical due to the extensive data requirements such processes entail.
(a) LLaMA 7B and LLaMA 13B.
(b) LLaMA 13B and LLaMA 33B.
Figure 10: Output cosine similarity Transformer layer groups of two models.
In BlockLLM, we leverage the output vocabulary probabilities as a means to ascertain such equivalence. By feeding the layers with the same data, we convert the output of each Transformer layer into vocabulary probabilities and compute the cosine similarity of the probability as an indicator of the degree of equivalence. In Figure 10, we have shown the output cosine similarity of Transformer Layer groups partitioned from different-sized models. The average similarity is 0.9841.
4.2 Model Partitioning
With the evaluated equivalence, BlockLLM incorporates an automated mechanism that breaks down LLMs into blocks. Each block then becomes an independent engine during serving.
Principles of constructing block zoo. BlockLLM follows two guiding principles: (1) avoid over-partitioning, (2) the preservation of architectural integrity. In this way, BlockLLM can reuse the blocks with minimal additional effort.
The rationale for reducing redundancy is to prevent unnecessary duplication. If one model component has no variant, it is meaningless to over-partition it into excessively small blocks. Such miniature partitioning is inefficient both in terms of computational power within each block and due to the increased overhead from data transfer between blocks.
The second principle directs BlockLLM to partition the model only at clear architectural boundaries. Based on PE and FF techniques we have adopted [37, 4, 25, 12], BlockLLM sets the following network components as the finest-grained network components that can compose a block: $attention$, $ffn$, $embedding$, $lm\_head$. This approach ensures that the output of one block seamlessly serves as the input to another, facilitating a streamlined chain of blocks for each request. BlockLLM avoids intricate arithmetic operations between blocks, simplifying the challenge of assembling them during the serving process. An illustrative example is the application of LoRA to two Linear operations within an Attention layer. Isolating LoRA as a separate block would lead to the partitioning of the Attention layer into five blocks, requiring BlockLLM to sum up the outputs of the LoRA block and the foundation Attention block. Therefore, in BlockLLM, each Attention with LoRA layer should be a separate block.
Figure 11: An example of BlockLLM’s lazy partitioning.
Lazy partitioning. BlockLLM employs a lazy partitioning strategy for all fine-tuned models utilizing a foundation model, slicing models as needed without preset restrictions on the size or architecture of the blocks. For PE models, it is partitioned into two blocks, the foundation model and its PE adapter.
As for FF models, after BlockLLM evaluates equivalence between model components ( 4.1), we perform lazy partitioning. As shown in Figure 11, for a FF model adapted from the foundation model (step 1), BlockLLM sets a threshold $T$ and group all connected network components (e.g., transformer layer 1 from the fine-tuned model and the foundation) that have equivalence exceeding $T$ into a block (step 2). This action results in 4 blocks, with two blocks being considered equivalent. When a newly added PE model (e.g., LoRA) arrives (step 3), BlockLLM (1) preserve its PE adapters (e.g., AttLoRA1, AttLoRA2), (2) find all existing blocks that share parameters (e.g., FFN1, FFN2) with the new model and (3) attempts to partition the existing blocks to enable block reuse. LoRA here has a different Attention layer. Therefore, we further partition the foundation model into four blocks so that they share the FFN layers in separately-provision blocks. This action results in 9 blocks with two chain-of-blocks being equivalent.
It is important to note that BlockLLM uses blocks as logical basic units of resource provisioning, and the granularity of blocks does not necessarily introduce overhead during serving time: BlockLLM considers overhead of pipelining blocks at fine-granularity and could choose to allocate consecutive chain of blocks in a single agent. The engine executes these blocks as a single network sub-component.
4.3 Stitching Block
Now that BlockLLM has partitioned LLMs into blocks, we tackle the third question. When two blocks of different embedding sizes are considered equivalent, such as Blocks B and D illustrated in Figure 9, how to route requests originally intended for, say Block B to D, becomes a unique challenge.
Stitching blocks. We propose to use a stitching block, a concept that has been recently explored [30]. We follow the principle of fine-tuning LLMs to obtain a stitching block. However, training a unique stitching block for each potential block pair is impractical. Thus, our objective is to design a generalizable stitching block. When multiple equivalences are identified between blocks originating from the same two foundation models, requests can be seamlessly redirected at any stitching point using the stitching block.
Stitching Block GPU hours LM Head Cosine Similarity
(4096, 5120) 4.33 0.9732
(5120, 4096) 4.84 0.9683
(4096, 8192) 6.32 0.9798
(5120, 8192) 5.85 0.9612
Table 3: Costs of training a stitching block in different sizes on A100 GPUs and output similarity with the large model.
We use a Linear layer to serve as the stitching block. This stitching block is trained while maintaining the parameters of other blocks static. To generalize the stitching block, we encode the positional information of the stitching point as an extra dimension. The position value is the sum of the positions of the head block and the tail block in its original foundation model. During training, it is initially placed at a shallow stitchable layer and progressively moved to deeper ones if there is any. This approach follows the idea of reusing layers in [22, 41]. Table 3 shows the training costs and the performance of the stitched model compared with the large model.
The underlying hypothesis of our approach is rooted in the Transformer-based architecture of LLMs: the output embedding from any Transformer layer can be interpreted as natural language. This characteristic presents the opportunity to engineer a stitching block with the unique capability of translating the output embedding from one dimension to another. This strategy allows the integration of varied block sizes.
5 Online Serving
We now present BlockLLM’s online serving system. We first explain how BlockLLM deals with stateful KV cache (C2,  5.1). Then, we discuss BlockLLM’s two strategies to mitigate the impact of communication overhead (C3, 5.2, 5.3).
5.1 KV Cache Coordination
Memory bandwidth-bound KV cache. Efficient stateful coordination of the KV cache is critical for auto-regressive LLMs in BlockLLM because memory bandwidth constraints on the KV cache emerge as a significant bottleneck, as evidenced by plenty of studies. Existing systems serve one batch of requests at a time and thus only consider the trade-off between recalculating the KV matrices and the cost of caching the KV matrices in the device memory. This trade-off reaches a point— determined by factors such as device type, model architecture, and request sequence length—where caching becomes more efficient than recalculation. However, as request sequences become longer, memory bandwidth constraints increasingly invariably become a performance-bounding factor when loading the KV cache [18].
I/O and recalculation cost. As aforementioned in 3.2, BlockLLM’s design complicates the problem. The assumption that requests will consistently be processed by the same block instances throughout all iterations no longer holds. To exploit the KV cache, the I/O cost of transferring cache between block instances may be unavoidable. We thus examine the cost implications within BlockLLM’s design space, considering two potential scenarios when a request batch currently residing on device $d_{i}$ initiates a new iteration.
In the first scenario, the request revisits the block instance on device $d_{j}$ that retains its KV cache. The latency is thus influenced by P2P request transfer overhead and the loading cost of KV cache. The increased latency can be expressed as:
$T_{transfer}^{w/kv}=\frac{D^{\prime}_{req}}{B_{net}(d_{i},d_{j})}+\frac{D_{cache}}{B_{mem}(d_{j})},$
where $D^{\prime}_{req}$ is the size of the newly-generated token and $D_{cache}$ is the size of its KV cache. $B_{net}(d_{i},d_{j})$ is the network bandwidth between two devices and $B_{mem}(d_{j})$ is the device memory bandwidth of $d_{j}$.
In the second scenario, the request is dispatched to a different block instance on device $k$ without the memory of its KV cache. The increased latency encompasses P2P request transfer overhead and either the P2P transfer cost of the KV cache or the recalculation cost of the KV cache. The resulting increased latency is formulated as:
$\begin{split}T_{transfer}^{w/kv}&=\min(\frac{D^{\prime}_{req}}{B_{net}(d_{i},d_{k})}+\frac{D_{cache}}{B_{net}(d_{j},d_{k})}+\frac{D_{cache}}{B_{mem}(d_{k})},\\ &\frac{D_{req}}{B_{net}(d_{i},d_{k})}+\frac{D_{cache}}{B_{comp}(d_{k})}),\end{split}$
where $D_{req}$ encapsulates the entire request, including the original input and all subsequently generated tokens. $B_{comp}(d_{k})$ denotes the computation capacity (i.e. FLOPS) on device $k$. We consider two cases for KV cache handling: recalculating the KV cache or transferring it. The former entails the transfer of both the request and its associated KV cache to the target block instance and the cost of loading the KV cache during computation, while the latter involves the transfer cost of the complete sequence for recalculation of the KV matrices.
Prioritize KV cache owner. Comparing these two scenarios, it becomes evident that transferring the KV cache to a new block instance is consistently less efficient than returning to the original KV cache owner. As for recalculating KV on the new block instance, it is also less efficient because caching typically outperforms recalculations, and network bandwidth is generally not as high-capacity as memory bandwidth. We follow this insight and perform best-effort coordination. As long as the candidate block instances have the same status (i.e., queuing time), BlockLLM’s agent should give priority to dispatch the request to the block instance that holds the associated KV cache. In 5.3, we discuss in detail how BlockLLM’s agent picks the suitable block instance.
Ownership of KV cache. BlockLLM is designed to store a KV cache for multiple requests, making it essential to distinguish the ownership of each KV cache matrix. Each agent maintains a dictionary mapping requests to the addresses of their corresponding KV caches. When processing a batch of requests, an agent of BlockLLM looks up this dictionary using the hashed request ID to retrieve the memory address of the relevant KV matrices. Additionally, multiple devices can hold a KV cache for a single request. To avoid this redundancy, the most recent copy is retained and BlockLLM’s scheduler periodically removes the others. When BlockLLM’s scheduler receives a completed request, it lets the agents possessing the KV cache of that request delete the associated memory.
5.2 Speculative Execution
Figure 12: Speculative execution of Block B.
Request transfer overhead hurts the overall throughput in two ways. First, it directly and inevitably prolongs inference latency due to its inherent blocking characteristic: each subsequent block’s operation is contingent upon receiving output from the preceding block, creating a strong dependency. Second, it exhausts network resources; each redirection of a request to a different server incurs network involvement. Moreover, as BlockLLM accommodates larger batch sizes than traditional serving systems, the size of data transferred increases, consequently consuming more resources. In BlockLLM, we propose to tackle the first with speculative execution and the second with block scheduling and placement.
Reduce latency via ahead-of-time prediction. Considerable efforts have been dedicated to enhancing P2P network transfers [33, 46], and such optimizations could be leveraged in BlockLLM to diminish inference latency. However, our approach seeks to offset any latency increases by accelerating the inference pipeline itself. Drawing from the design in the operating system, we exploit speculative execution to expedite the inference process. Instead of awaiting actual outputs from a preceding block, it predicts the outcome, allowing subsequent blocks to proceed based on these predictions. Should the predictions align with the actual results, this speculation can lead to reduced latency. As illustrated in Figure 12, when Block B adopts speculative execution using Pred B, the inference latency is shorter if it is verified to be a correct prediction.
The challenge of a successful speculation lies in both the time and the accuracy of the prediction. If the prediction takes a longer time than the actual computation, the benefits of speculation are nullified. If the accuracy of predictions is low, reliance on the actual output becomes inevitable to ensure computational correctness, rendering the speculative efforts both redundant and a potential drain on resources.
Block Pruned Percentage Cosine Similarity Computation Speedup
5th Attention 49.71% 0.7 37.43$\times$
5th FFN 49.83% 0.77 1.71$\times$
5th Transformer 49.79% 0.94 22.91$\times$
5th–7th Transformers 49.79% 0.84 18.55$\times$
Table 4: Performance of blocks’ surrogates in terms of prediction similarity and computation time. Blocks are pruned from LLaMA 7B based on the structured pruning technique described in [23].
Block surrogates. We leverage insights from existing work and empirical studies, coupled with our block zoo, to construct high-fidelity surrogates for the blocks. The surrogates are used for prediction. First, we consider pruned models. One typical pruning technique is sparse models. While sparsity typically results in a reduction of computation FLOPs, it does not always correspond to a commensurate decrease in computation time. Hence, our focus is on pruning techniques that meaningfully cut down computation time. We implement a strategy inspired by existing literature [23] to construct surrogates for our blocks. In particular, we perform structured pruning by selectively removing sub-structure that generates less critical impact towards the model output and then add fine-tuned LoRA for performance recovery after the pruning process. Table 4 shows the output similarity and computation time of four blocks and their surrogates. Second, we exploit equivalent blocks with different computation costs in BlockLLM’s block zoo but in different computation costs in BlockLLM’s zoo. For instance, a block partitioned from a 7B LLaMA model can serve as an effective surrogate for a block from a 30B LLaMA model, which is adopted by [26].
Figure 13: Possible scenarios when speculative execution is consecutively applied to all blocks.
Bottleneck-pinpointing speculation. Naively application of speculative execution to blocks is not a guarantee of latency improvement due to the necessity of verifying predicted results. In Figure 13, consider an inference pipeline composed of three blocks where speculative execution is applied to each. Ideally, if all speculations are destined to be correct, the inference latency is expected to be reduced from T6 to T2 ( 1). However, consecutive prediction verification can diminish these gains, especially when erroneous predictions happen at the latter part of the chain. For instance, if a prediction error arises at Block C (last block in the chain) ( 4), the latency reduction would only be 0, compared to a reduction of T6-T3 if the error occurred earlier at Block A 2. Therefore, the optimal inference ( 5) one could achieve is T5 instead of T2. In the worst case, where all three speculations are incorrect, there would be no latency reduction at all. Practically, due to the concurrent execution of blocks and their surrogates, this could even result in additional slowdowns.
To avoid such outcomes, we follow two rules when implementing speculation. First, we avoid enabling speculation on consecutive blocks to minimize accumulated errors and wasted resources. Second, speculation is not applied at the last block in the chain because the final output, once relayed back to BlockLLM’s scheduler, is not correctable. Thus, during serving, we selectively apply speculation to the top-$k$ bottleneck blocks—those characterized by their computation intensity or frequent use.
5.3 Request and Block Scheduling
We then explain BlockLLM’s scheduling for requests and blocks.
Adaptive serving. BlockLLM’s scheduler and agents perform adaptive serving when dispatching the request to a suitable block instance. With the help of block zoo, the candidate block instances for each request include the block that is explicitly listed in the chain and its equivalent blocks.
BlockLLM’s request scheduling strategy is to prioritize the KV cache owner instance as long as its device memory can hold the request data; otherwise estimate the latency of each candidate instance and schedule the request onto the instance with the lowest latency increase. When the KV cache owner is not available, the components of the latency estimation of a candidate block instance residing on device $c$ are as follows:
$\begin{split}Latency_{d_{c}}&=T_{queue}+T_{compute}+T_{transfer}+T_{load},\\ T_{queue}&=\sum_{i=1}^{n}{Comp(req_{i})},\\ T_{compute}&=Comp(req),\\ T_{transfer}&=\begin{cases}T_{transfer}^{w/kv}&\text{agents dispatch,}\\ \dfrac{D_{req}}{B_{net}(s,d_{c})}&\text{scheduler dispatches,}\end{cases}\\ T_{load}&=\begin{cases}0&\text{idle device,}\\ \dfrac{D_{b^{\prime}}}{B_{mem}(d_{c})}+\dfrac{D_{b}}{B_{net}(d_{c})}&\text{loading overhead.}\\ \end{cases}\end{split}$
We incorporate four key factors: queuing time, computation time, transfer time, and the overhead associated with block switching. Queuing time accounts for the duration required to process all $n$ queuing request batches. Transfer time is subject to who initiates the dispatch: if done by the scheduler $s$, the request batch is at the start of the chain and no KV cache is involved. Otherwise, agent dispatching includes the costs of transferring the request and its KV cache (discussed in 5.1). The overhead from switching block depends on the current status of the device; if idle, the block’s loading can be overlapped with other operations, thus zero overhead. Otherwise, this loading overhead includes the time to move out the existing block $b^{\prime}$ and load the new block $b$.
Block resource allocation. BlockLLM’s scheduler determines the resource allocation of blocks, similar to a traditional serving system. We discuss BlockLLM’s strategy for enabling independent per-block scaling and speculative execution. For scaling, we follow a straightforward policy centered on the queue length of the block instance. If the queue length exceeds tof the maximum queue length (i.e. saturating all the device memory), we scale onto more devices starting from the most heavy-loaded block instances. If the block instance to be scaled possesses requests’ KV cache, we balance the load by moving the state as well. As for speculative execution, BlockLLM actively applies to the top-$k$ time-consuming block instances to accelerate the inference pipeline. BlockLLM deploys surrogates using a dedicated stream on the same device where the blocks to be speculated are situated.
Locality-aware block placement. BlockLLM’s second effort to mitigate the transfer overhead is the strategic placement of blocks, aiming to reduce the reliance on network resources. During placement, BlockLLM prioritizes locality—ensuring that blocks with frequent direct inter-dependencies are close to each other. Ideally, these blocks are placed on the same server. Such an arrangement leverages the full potential of high-capacity intra-server connections, such as NVLink interconnects, for the transfer of requests and KV cache, rather than resorting to the constrained inter-server links.
In BlockLLM, the locality is quantified by monitoring historical traffic, specifically by a counter recording the frequency of being directly inter-dependent between two blocks. Block pairs with a higher locality value are placed within the same server. Furthermore, BlockLLM’s scheduler dynamically adapts to evolving traffic patterns; should the observed locality change, block instances are migrated as necessary to align with the new pattern.
6 Implementation
We have implemented a prototype of BlockLLM. It is compatible with HuggingFace models. We use NCCL as the communication backend to transfer data among servers.
Profiling. To support the online serving system, BlockLLM conducts profiling on blocks. First, we profile the computation time of each block. It includes the computation time measured under different batch sizes. This also applies to the surrogates of each block and the performance when the block and its surrogates are multiplexing. Second, we obtain the communication time from one device to another device in different transfer sizes using NCCL primitives. Third, we measure the overhead of loading the block engine from disk to host memory and to device memory and vice versa.
Batching. The computation efficiency improves with a larger batch size. However, forcing the blocks to adopt a fixed large batch size would complicate the problem because it involves reorganizing the requests from different batches/applications. Therefore, we loosely encourage batching within each block instance. Upon receiving a new batch of requests, BlockLLM’s agent puts it into the queue if there are any and attempts to pack it with its direct neighbors as long as the combined batch does not exceed the upper limits of the batch size of the block. If there is no queuing request, BlockLLM’s agent sends the request batch directly to the instance for processing. Additionally, in case one request in the batch reaches EOS at the last chain of block, it is removed from the batch and relayed to the scheduler for response.
Request dispatching. BlockLLM’s agents adopt a FIFO+priority queue, where priority is assigned to requests that have left KV cache memory and may revisit in the future. Each block instance maintains a countdown clock associated with every auto-regressive request based on the time estimation of one iteration, anticipating its potential return. As long as the countdown remains active, BlockLLM’s agent prioritizes the processing of returning requests.
The scheduler and agents dispatch the request in slightly different ways. BlockLLM’s agent needs to (1) determine the candidate block for each request and pack them in different buckets and (2) obtain the placement of the block candidates. BlockLLM’s agent initiates a broadcasting request, and those available agents hosting the candidates would respond. BlockLLM’s scheduler, on the other hand, keeps a live record of the placement of every block instance in the cluster, thus easing the dispatching process.
7 Evaluation
We evaluate BlockLLM using testbed experiments. We seek to answer two main questions:
  • Does BlockLLM improve the overall throughput and resource utilization compared to existing solutions? (7.2)
  • For online serving, how much latency reduction is contributed by BlockLLM’s key designs? (7.3)
7.1 Setup
Cluster. Our cluster has four servers. Two are equipped with two A100 GPUs with 80GB memory, and the other two have four A100 GPUs with 80GB memory. The servers are interconnected with 100Gbps network.
LLM Models. We use two foundation models in different sizes: 6B Chat-GLM [44], 7B LLaMA, and 13B LLaMA [39]. We consider fine-tuning techniques, including FF (Vicuna [4]) and PE (Prefix-Tuning, Adapter, BitFit, and LoRA). We have 20 fined-tuned LLM models, each representing an application.
Workload. We generate a synthetic workload trace (Figure 15) to evaluate BlockLLM’s performance. We use uniform distribution to generate a mean rate for each application. The mean rate is used as the weight to calculate the number of requests for each application, so some applications are more popular. Then, the arrival time of each request in one application is generated using a Poisson process with the given mean rate. The trace is 20 minutes, and there are 400 requests. The generation approach is similar to [34].
Figure 14: 20-min workload.
Figure 15: Latency CDF.
Baselines. We benchmark BlockLLM with two other solutions.
  • Per-Model provisioning (PM): Each LLM is deployed independently. Models are scaled along with traffic demand.
  • Parameter Sharing (PS): We merge PE LLMs from the same foundation and deploy them as a complete unit, similar to S-LoRA [34]. Since we support various fine-tuned applications, we use branching to differentiate each application.
Metrics. We evaluate BlockLLM with metrics, including latency of serving one request, throughput (number of tokens per second), communication costs (percentage of time spent on data transferring), GPU utilization (SM efficiency), and device memory consumption.
BlockLLM’s configuration. In this evaluation, we apply speculation to the top 10% bottlenecked block instances sorted by the time of completing their request queues. BlockLLM’s scheduler checks redundant KV cache every one minute. We set 0.95 as the cosine similarity threshold for determining whether BlockLLM’s surrogate prediction is accurate and 0.98 as the threshold for determining equivalence. The maximum sequence length allowed is 1024. BlockLLM trains the stitching block with the MMLU [14].
7.2 Overall Performance
We first compare the overall performance of BlockLLM. The default number of applications is 20.
Figure 16: Throughput and GPU utilization.
Figure 17: GPU utilization change over time.
Latency and throughput. Figure 15 depicts the CDF of the latency of completing a request in BlockLLM. BlockLLM’s median latency is 3.34, comparable with PM (3.39) and PS (3.87) provisioning. The reduction of 95%ile latency is more significant, 33.5% and 23.4% compared with PM and PS. The throughput of BlockLLM is 1.71x of PM and 1.46x of PS provisioning (Figure 17). Since BlockLLM allows larger batch sizes with higher efficiency, the tail latency is improved significantly. Though PS can serve various applications within one model, the cost of one inference process grows due to branching.
Figure 18: Memory usage of parameters and request-related data.
Figure 19: Performance changes when the the number of applications grows.
GPU utilization. We also measure overall GPU utilization indicated by SM efficiency and memory consumption in the cluster. We monitor the end-to-end serving process, and the average GPU utilization is improved by 20.1% and 4.8% compared with PM and PS provisioning. Figure 17 depicts the GPU utilization change over the serving process. With the help of adaptive serving, BlockLLM efficiently dispatches requests under the existing deployment status to avoid model frequent model loading and unloading. As for memory consumption, we measure the memory consumption of model parameters and request-related matrices (including input, intermediate activations, output, and KV cache). In the best case, BlockLLM takes 16.1% less space on model parameters and 24.1% more space on request-related matrices, indicating more data are being processed.
Number of applications. We increase the number of applications from 10 to 30 with the same approach in 7.1. When the number of applications grows to 30, and resources become increasingly constrained, BlockLLM’s performance gain is more prominent, achieving a 37.4% reduction in 95%ile latency and a 1.85x improvement in overall throughput (Figure 19). BlockLLM exploits block reuse to serve more applications under the same resource constraints and its flexibility brings efficiency in request serving.
7.3 Effectiveness of BlockLLM’s Design
We then evaluate the effectiveness of BlockLLM’s key designs using an ablation study.
Figure 20: Cosine similarity of requests using adaptive serving and predetermined block chains.
Figure 21: Ablation study of KV cache coordination strategy.
Figure 22: Ablation study of speculation: no speculation and perfect speculation.
Figure 23: BlockLLM compared with fragmentation-minimized placement.
Adaptive serving. We summarize the statistics of adaptive serving in BlockLLM. When all the requests allow adaptive servings, 136 out of 400 requests are served with adaptive chains of blocks. We compare the output vocabulary probabilities of these requests to the ones without adaptive serving and depict the cosine similarity CDF in Figure 21. The average similarity is 0.88. We disable adaptive serving in BlockLLM and the 95%ile latency and throughput are degraded by 15.6% and 23.7%, respectively.
Best-effort KV cache coordination. BlockLLM performs best-effort dispatching by prioritizing the device with its KV cache memory. We consider two other solutions to verify if BlockLLM’s strategy is efficient: (1) All the KV cache are obtained using recalculation. (2) We always route the request back to the least busy device and let the KV cache owner transfer the cache to the instance. Figure 21 shows the median and 95%ile latency, throughput, and communication costs normalized to BlockLLM. The 95%ile latency is increased by 1.23x using recalculation and by 1.36x using least-busy-device routing. The communication costs are reduced significantly to 0.36 of BlockLLM when using recalculation and increased to 1.28x when using least-busy-device routing.
Speculative execution. We evaluate the effectiveness of BlockLLM’s speculative execution via two approaches in Figure 23. First, we disable speculative execution and compare BlockLLM’s performance change. The median and 95%ile latency inflates by 11.3% and 31.6%, compared with BlockLLM with speculative execution. In this evaluation, speculative execution has been performed 231 times, 192 out of which have made predictions accurate enough to avoid correction. Second, we replace our surrogates with pseudo surrogates, where we assume their predictions are always accurate with an ideal 1/50 computation time of the speculated block. The ideal 95%ile latency and throughput are 87.3% and 112.8% of BlockLLM with real surrogates. This shows that our surrogates are both efficient and accurate.
Locality-aware block placement. We compare the communication costs between BlockLLM’s locality-aware placement and the widely adopted fragmentation-minimized placement. Figure 23 shows the average performance change of using fragmentation-minimized placement. The median and 95%ile latency is increased by 12.6% and 18.2%. The communication costs of one request sum up all the transfer costs, therefore presenting a significant inflation of 73.4%. The locality-aware placement has reduced 72.3% inter-server communications compared with the fragmentation-minimized placement strategy.
8 Discussion
Stitching models in different sizes. The concept of model concatenation, particularly across varying sizes, has been investigated in prior research with notable efforts in Convolutional Neural Networks and Transformers [10, 30]. In BlockLLM, we introduce a simple technique to validate the potential of stitching blocks of disparate sizes, ensuring that such integration is only executed between blocks verified to be equivalent. However, it has limitations; it typically requires retraining and the maintenance of the original layer order. To overcome these limitations, a more advanced strategy is imperative based on a deeper comprehension of the models’ underlying representations and functionalities.
Other opportunities. BlockLLM’s design brings other practical benefits. (1) Partial updates. Parameter updates can be applied in block-granularity as well. With actor model [15, 24], updates can be performed concurrently with the serving process without interruption. Given the update frequency difference between the foundation and fine-tuning models, partial updates could effectively upgrade the system whenever necessary. (2) Early-exit [38] is a promising approach to reducing computation costs. However, batching limits its potential because the requests requiring the deepest computation always bottleneck the process. BlockLLM’s design choices naturally fit with early-exit, where requests can jump out of the batch at any block, which is similar to DVABatch [7], which tries to enable early-exit via multi-entry multi-exit batching.
9 Related Work
System for LLM models. Model-internal optimization including FasterTransformer [2], PagedAttention [18], FlashAttention [8] and FlexGen [35] are complimentary to BlockLLM’s contributions. Orca [42] introduces iteration-level scheduling to handle the sequence length difference among multiple requests in the batch. SpotServe [27] realizes LLM serving on preemptible instances with communication cost-minimized migration plan and stateful recovery. ORCA and SpotServe can be integrated with BlockLLM’s scheduler and agent. PetS [45] expresses four representative PE techniques in a unified representation so that the requests from four applications can be executed simultaneously, similar to Parameter Sharing evaluated in 7.
Auto-regressive. The inherent auto-regressive nature of LLMs limits their serving efficiency, making parallel decoding a primary focus of recent research. SpecInfer [26] and [19] circumvents this constraint through speculative execution. It employs smaller LLM models to predict multiple tokens in advance. [29, 11] further improve the solution by removing the need for surrogate models. Instead, they use the existing model together with a few look-ahead tokens to make predictions. All these strategies aim to generate multiple preliminary drafts, thereby facilitating earlier production of the subsequent token. BlockLLM, however, implements speculation with a more refined approach. BlockLLM applies speculative execution at the block-level in a selective manner, optimizing the generation process at the individual token level.
10 Conclusion
We present BlockLLM, a multi-tenant finer-grained serving system tailored for LLM workloads. In BlockLLM, we show the effectiveness of improving throughput by allowing model component reuse with blocks. We enable adaptive serving, effectively coordinate multiple requests’ KV cache, and mitigate the communication costs to improve serving efficiency. Experiments show the efficiency of BlockLLM. We plan to extend BlockLLM to support more models and evaluate it on larger-scale clusters.